OpenStudy (anonymous):

solve the equation (x+2) ^2 -6 = 19

OpenStudy (anonymous):

\[(x+2)^{2}=19+6\]\[(x+2)^{2}=24\]\[(x+2)=\sqrt{24}\]\[x=\sqrt{24}-2\]